OpenStudy (anonymous):

difference quotient f(x)-f(a)/x-a f(x)=-x^2 a=1

myininaya (myininaya):

\[\frac{f(x)-f(a)}{x-a}\] \[f(x)=-x^2 => f(a)=-a^2\] Plug in and simplify.